OpenStudy (anonymous):

I need help evaluating the following: (equation is below) thank you. Step by step instruction would be appreciated, so I can learn and keep in drilled in my mind! I need help evaluating the following: (equation is below) thank you. Step by step instruction would be appreciated, so I can learn and keep in drilled in my mind! @Mathematics

OpenStudy (anonymous):

\[(7x^2+8x+6)/(\sqrt(x))\] I know its quotient rule, but the radical kind of throws me off!

OpenStudy (anonymous):

hey u know Sqrt(x) = x^(1/2)

OpenStudy (anonymous):

\[(7x^{2} + 8x +6)/(\sqrt{x})\] \[(7x^{2}+8x+6)/(x ^{1/2})\]

OpenStudy (anonymous):

nice one sheg

OpenStudy (anonymous):

now divide it !

OpenStudy (anonymous):

\[(7x^{2} + 8x + 6) X ( x^{-1/2})\] \[(7x^{2-1/2} + 8x^{1-1/2} + 6x^{-1/2})\]

OpenStudy (anonymous):

\[7x^{3/2} +8x^{1/2} + 6x^{-1/2}\] this is your solution

OpenStudy (anonymous):

You mention quotient rule. Do you need to find derivative?

OpenStudy (anonymous):

yeah, sorry i didn't mention that clearly

OpenStudy (anonymous):

was this u looking for

OpenStudy (anonymous):

In this case it's very easy for 7x^(3/2) +8x^(1/2) +6x^ (−1/2) . You just do simple derivative for each term (x^n)'=n*x^(n-1)

OpenStudy (anonymous):

I need the derivative of the function at the top using the quotient rule

OpenStudy (anonymous):

thank you both sheg and Marina!

OpenStudy (anonymous):

you are welcome. You can do it both ways: 1) using quotient rule 2) how I advise you. Your final answer will be (21x^2+8x-6) / 2x^(3/2). I think the second way is easier, but you can make your own conclusion by doing both. Have fun!

OpenStudy (anonymous):

^ how did you get that derivative ... I don't follow , sorry!